OpenStudy (anonymous):

how many triples of natural numbers have sum of 11

OpenStudy (anonymous):

is it 13!/(11!*2!) -12!*2!/(11!*2!)= 78-12=66

OpenStudy (anonymous):

66

OpenStudy (anonymous):

is it 13!/(11!*2!) -12!*2!/(11!*2!)= 78-12=66

OpenStudy (raden):

because the numbers must be natural numbers, so the sum of triples are : (11-1)C(3-1) = 10C2 = 45

OpenStudy (anonymous):

u can re-write in (11+3-2)C(3-1) =66

OpenStudy (raden):

i have tried by manual (exploration method) too, it can be 45, @matricked if im not mistake

OpenStudy (anonymous):

@RadEn can u explain...

OpenStudy (raden):

i want to answering by exploration let the numbers are a,b, and c (with a,b, and c are natural numbers) so, we get an equation a+b+c=11 * case I, if a=1 then b+c=10 or c=10-b to get c is natural number, so the value of b must be = 1,2,3,..,9 therefore there are 9 order pairs (a,b,c) * case II, if a=2 then b+c=9 or c=9-b to get c is natural number, so the value of b must be = 1,2,3,..,8 therefore there are 8 order pairs (a,b,c) .... so on, until : * case 9, if a=9 then b+c=2 just 1 option only, it satisfied for b=c=1 therefore there is 1 order pair (a,b,c) so, the total order pairs (a,b,c) = 9+8+7+6+...+2+1 = 45

OpenStudy (raden):

am i right, @matricked ?

OpenStudy (anonymous):

@RadEn yup dear you are absolutely right.. i have aslo verified the same by writing a simple program and found out to be 45 i have also include the ways in which a=0 (not a natural number ) and c=0 (not a natural number and hence it will be 66 -(10+10+1)=45 ways ur explaination is also very clear thank you very much for correcting me @choze my ans isn't needs correction so follwo RadEn's method...